OpenStudy (anonymous):

The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 a.m. and the airplane starts from the same place at 11 a.m., they will have traveled the same number of miles by 1:00 p.m. If r represents the rate of the automobile, how many miles does the plane travel?

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

Let the average speed of the airplane be x. x=200+r speed=distance/time distance=speed*time Time taken by airplane=13.00-11.00=2 Time taken by automobile=13.00-3.00=10 Distances are equal \[\frac{ 200+r }{ 2}=\frac{ r }{ 10 }\] Can u find r @david17?

OpenStudy (anonymous):

t* r = d Solve the following for r:\[10 r=2(200+r)\]r = 50\[10*50 = 500 \]

OpenStudy (ajprincess):

oops sorry. \(2(200+r)=10*r\)

OpenStudy (anonymous):

@ajprincess Thank you for the medal.

OpenStudy (anonymous):

200r 2(200 + r) 10r @ajprincess is a

OpenStudy (ajprincess):

Welcome @robtobey:) nd I didnt get ur question @david17.